Please use a calculator for the following problems and please show work

1) Brianne plans to deposit $100,000 today into a fund that will be needed at the end of 6 years. She will receive 12% interest on the fund balance. What is the fund balance at the end of year 6 assuming semiannually compounding

2) What is the present value of $10,000 to be received each year for 20 years if the payments start immediately. Assume a discount rate of 4%

3) Robert wants to withdraw $100,000 a year for 30 years after he retires. He will probably work for 30 years and earn 4% interest during the whole time. How much does Robert have to contribute to her pension fund each year so she can retire on $100,000 a year
Answer #1

The answer has been presented in the supporting sheet. All the parts has been solved with detailed explanation and format. For detailed answer refer to the supporting sheet.

2 3 4 5 Answer Part 1) Future value = amount invested*(1+i)an = 100000*(1+(12%/2)^(6*2) = $201219.65 (rounded off to 2 decima